OpenStudy (tylerd):

an ellipse that has a center at (9,2) a focus at (6 , 2) , and a vertex at (14 , 2 )

OpenStudy (tylerd):

im trying to solve for b if \[\frac{ (x-c)^2 }{ a^2 }+\frac{ (y-d)^2 }{ b^2 }=1\]

OpenStudy (tylerd):

a=5, c=9, d=2

OpenStudy (tylerd):

but how to find b?

OpenStudy (tylerd):

@jim_thompson5910